# Exposing the Cracks: Vulnerabilities of Retrieval-Augmented LLM-based Machine Translation

### v1

@@ Introduction
Given that REAL-MT consistently adopts retrieved context even when it is incorrect, we investigate the model’s confidence in these error-prone outputs. Specifically, under highly misleading noise ($\mathcal{N}_{\text{opposite}}$), does the model assign low confidence to its context-induced translations, signaling awareness of potential error? Such uncertainty would indicate good calibration and enable confidence-based self-correction or rejection mechanisms. Conversely, high confidence in erroneous outputs would reveal a dangerous overtrust in retrieval, undermining system trustworthiness. To assess this, we measure confidence using the entropy of the output probability distribution, with lower entropy indicating higher confidence. To ensure we evaluate the correct segment, we identify the target tokens corresponding to the source idiom through attention-based alignment: for each generated token, we compute its attention weights over the input and determine if it attends primarily to the idiom or contextual cue. The longest continuous sequence of idiom-aligned tokens is treated as the translated idiom span, and we report average entropy over this span.
As shown in Figure 4, the model exhibits increased confidence in the translation of idiom spans under the $\mathcal{N}_{\text{opposite}}$ setting, even though fidelity drops significantly. This inverse relationship reveals a critical failure in calibration: the model becomes more certain of its outputs despite their inaccuracy, suggesting that it implicitly treats the noisy context as authoritative and integrates it without sufficient verification.
Figure 4: Idiom-span translation confidence and Fidelity in Qwen2.5-7B-Instruct when adopting $\mathcal{N}_{\text{opposite}}$ context.
@@
To counter REAL-MT’s blind trust in noisy retrieval and its lack of self-verification, we investigate training-free strategies for on-the-fly rejection of unreliable contexts and training-based methods to teach models to discern and downweight misleading knowledge.
@@
In RAG scenarios, prior work has explored the fusion of external knowledge to handle conflicts with internal knowledge. Given the similarity to REAL-MT, where noisy retrieved contexts may mislead translation, we investigate CK-PLUG (Bi et al. 2025), a training-free method that dynamically controls knowledge reliance based on context reliability. CK-PLUG computes Confidence Gain (CG) to measure the change in token-level entropy. For tokens with positive CG, which indicates the context is beneficial, the method blends the context-aware and internal distributions using $\alpha=0.5$. For tokens with negative CG, where the context may be harmful, external knowledge is fully suppressed. Although our analysis shows that REAL-MT exhibits poor calibration under noise, CG relies on relative confidence shifts rather than absolute certainty, potentially capturing whether the context stabilizes or disrupts the model’s internal prediction. We therefore evaluate CK-PLUG as a representative training-free strategy to test whether such entropy-based signals can still enable on-the-fly rejection of harmful contexts, despite overall miscalibration.
@@
To address REAL-MT’s tendency to blindly trust retrieved context even when it is misleading, we investigate training-based strategies that explicitly expose models to adversarial retrieval conditions during fine-tuning. Specifically, we construct training instances where the retrieved context conveys the opposite meaning of the source idiom, while the target translation remains correct. This encourages the model to learn to disregard misleading external knowledge and rely more on its internal representation when the two conflict. To prevent overcorrection (i.e., ignoring all context, including accurate ones), we also include noise-free conditions, No Context ($\mathcal{C}_{\text{none}}$) and Gold Meaning ($\mathcal{G}$) settings, during training. Our goal is to assess whether this fine-tuning scheme improves robustness to noisy contexts without sacrificing performance when retrieval is accurate.
@@
This design directly addresses REAL-MT’s core failure mode: blind trust in retrieved context. By exposing models to misleading contexts with correct supervision (ALI), we encourage them to learn that not all retrieved knowledge is reliable. CDA further teaches discriminative reliance by promoting trust in accurate context ($\mathcal{G}$) and rejection of adversarial noise ($\mathcal{C}_{\text{none}}$), a crucial step toward self-verifying integration. Including clean conditions prevents overcorrection and preserves the performance when retrieval is accurate.
@@
Both training-free and training-based strategies improve robustness but degrade performance under $\mathcal{G}$, with the trade-off most pronounced in low-resource settings. This tension arises because neither approach can dynamically modulate reliance on retrieved context; instead, they apply fixed or static policies that inevitably sacrifice either noise resilience or clean-context utility. The strong dependence on hyperparameters like the blending weight $\alpha$ and training data proportions further underscores their brittleness. These findings reinforce the need for self-verifying integration mechanisms that can reliably assess context reliability at inference time, rejecting noise while preserving the benefits of accurate knowledge.
Conclusions
In this paper, we address the critical gap in understanding REAL-MT’s reliability under noisy retrieval, a common yet overlooked challenge in real-world deployment. To this end, we propose a controlled noise synthesis framework and two new metrics, Fidelity and Context Adoption Rate (CAR), to systematically evaluate REAL-MT’s robustness. Our evaluation, using Qwen-series models across high-, medium-, and low-resource language pairs, reveals that low-resource pairs are most affected by noise due to their stronger reliance on retrieval, often leading to nonsensical translations. Surprisingly, large reasoning models (LRMs) with enhanced reasoning capabilities not only fail to mitigate errors but are more vulnerable, frequently rationalizing incorrect contexts. This stems from a dual failure: attention shifts away from source idioms to noisy content, combined with rising confidence despite declining accuracy, highlighting poor calibration and weak self-verification. We explore both training-free and fine-tuning strategies to enhance robustness, but these come with trade-offs, such as reduced performance in clean contexts. Our findings highlight the limitations of current REAL-MT systems and emphasize the urgent need for self-verifying mechanisms to critically assess retrieved knowledge, allowing models to filter out noise while preserving the benefits of accurate external information.

### v2

@@ Introduction
Given REAL-MT’s uncritical adoption of noisy context, we probe the calibration of its context-induced translation. Under $\mathcal{N}_{\text{opposite}}$ noise, does the model signal awareness of its mistakes through low confidence, or does it remain highly confident in its errors, indicating poor calibration? We quantify confidence as the entropy over the output distribution, focusing on the idiom span identified via attention-based alignment. To ensure we evaluate the correct segment, we identify the target tokens corresponding to the source idiom through attention-based alignment: for each generated token, we compute its attention weights over the input and determine if it attends primarily to the idiom or contextual cue. The longest continuous sequence of idiom-aligned tokens is treated as the translated idiom span, and we report average entropy over this span.
As shown in Figure 4, the model exhibits increased confidence in the translation of idiom spans under the $\mathcal{N}_{\text{opposite}}$ setting, even though fidelity drops significantly. This inverse relationship reveals a critical failure in calibration: the model becomes more certain of its outputs despite their inaccuracy, treating noisy context as authoritative. This highlights the model’s limitations in recognizing internal and external knowledge (Cheang et al. 2023, 2025).
Figure 4: Comparison of confidence and Fidelity with $\mathcal{C}_{\text{none}}$ and $\mathcal{N}_{\text{opposite}}$ for idiom-span translations in Qwen2.5-7B-Instruct. Lower entropy implies higher confidence.
@@
We investigate training-free and training-based methods to improve REAL-MT’s robustness to noisy contexts.
@@
In RAG scenarios, prior work has explored the fusion of external knowledge to handle conflicts with internal knowledge. Given the similarity to REAL-MT, where noisy retrieved contexts may mislead translation, we investigate CK-PLUG (Bi et al. 2025), a training-free method that dynamically controls knowledge reliance based on context reliability. CK-PLUG computes Confidence Gain (CG) to measure the change in token-level entropy. For tokens with positive CG, which indicates the context is beneficial, the method blends the context-aware and internal distributions using $\alpha=0.5$. For tokens with negative CG, where the context may be harmful, external knowledge is fully suppressed. We apply CK-PLUG to assess its effectiveness in improving robustness against noisy contexts in REAL-MT without requiring additional training.
@@
We investigate training-based methods to teach models to discern and reject misleading knowledge. Specifically, we construct training instances where the retrieved context is $\mathcal{N}_{\text{opposite}}$, while the target translation remains correct. This encourages the model to learn to disregard misleading external knowledge and rely more on its internal knowledge when they conflict. Our goal is to assess whether this fine-tuning scheme improves robustness to noisy contexts.
@@
ALI is designed to train models to discern and reject noisy texts. CDA further teaches the model discriminative reliance by enabling it to trust $\mathcal{G}$ and reject $\mathcal{N}_{\text{opposite}}$. We include $\mathcal{C}_{\text{none}}$ to maintain the model’s core capabilities in this condition.
@@
ALI achieves the highest Fidelity and lowest CAR under $\mathcal{N}_{\text{opposite}}$ (e.g., Fr→En: Fidelity 1.8, CAR 11%; Hi→En: Fidelity 0.8, CAR 26%), substantially outperforming CK-PLUG. Crucially, despite being trained only on opposite-meaning noise, ALI generalizes to other semantic distortions ($\mathcal{N}_{\text{literal}}$, $\mathcal{N}_{\text{semantic}}$), demonstrating an emerging ability to discriminate reliable from misleading knowledge. However, it shows no improvement under syntactic perturbations ($\mathcal{N}_{\text{struct}}$), indicating that its robustness is limited to semantic noise and does not generalize to syntactic distortions.
Both training-free and training-based strategies improve robustness but degrade performance under $\mathcal{G}$, with the trade-off most pronounced in low-resource settings. This tension arises because neither approach can dynamically modulate reliance on retrieved context; instead, they apply static policies that inevitably sacrifice either noise resilience or accurate context utility. The strong dependence on hyperparameters like the blending weight $\alpha$ and training data proportions further underscores their brittleness.
Conclusions
In this paper, we address the critical gap in understanding REAL-MT’s reliability under noisy retrieval, a common yet overlooked challenge in real-world deployment. To this end, we propose a controlled noise synthesis framework and two new metrics, Fidelity and Context Adoption Rate (CAR), to systematically evaluate REAL-MT’s robustness in knowledge-intensive translation tasks. Our evaluation, using both open- and closed-sourced models across high-, medium-, and low-resource language pairs, reveals that all models exhibit a high sensitivity to semantic noise. This is especially pronounced in low-resource scenarios, where models are more likely to uncritically rely on the noisy external context, often leading to nonsensical translations. Our experiments on LRM uncover that it is more vulnerable to noisy context than the non-reasoning version and, critically, tend to rationalize the noise even when aware of it, failing to demonstrate truthful reasoning. An analysis of attention mechanisms and model uncertainty reveals two failures: a shift of attention away from source idioms towards noisy context, compounded by a rise in model confidence despite a decline in accuracy. This points to critically weak self-verification and poor calibration. While both training-free and fine-tuning strategies can enhance robustness under noisy contexts, they do so at the cost of performance in accurate contexts. This trade-off underscores the necessity of developing models with self-reflective capabilities.
Acknowledgments
This work was supported in part by the Science and Technology Development Fund of Macau SAR (Grant Nos. FDCT/0007/2024/AKP, FDCT/0070/2022/AMJ, FDCT/060/2022/AFJ), the National Natural Science Foundation of China (Grant Nos. 62261160648, 62266013), the China Strategic Scientific and Technological Innovation Cooperation Project (Grant No. 2022YFE0204900), and the UM and UMDF (Grant Nos. MYRG-GRG2025-00236-FST, MYRG-GRG2023-00006-FST-UMDF, MYRG-GRG2024-00165-FST-UMDF, EF2024-00185-FST, EF2023-00151-FST).
A Prompt Templates of Controlled Noise Context Generation
A.1 Gold Meaning Extraction
Provide the meaning of the idiom “{idiom_in_source}” as used in the source sentence, based on the context of the sentence and its translation. Source Sentence: {source} Translation: {target} Respond only with the idiom’s meaning in English without any explanation, formatted as a JSON object: { “output”: “meaning content” }
A.2 Structure-Perturbed Gold Meaning Generation
Goal: Create text with SUBTLE GRAMMATICAL ERRORS while preserving the core meaning of the input. Input: {groundtruth_meaning} Methods (MUST apply at least one): 1. Swap adjacent phrases/clauses to create awkward phrasing (e.g., “New problems that will arise” → “Problems new that will arise”) 2. Move modifier to incorrect position (e.g., “The quickly running child” → “The running quickly child”) 3. Slight word order inversion in short phrases (e.g., “good ideas” → “ideas good”) Constraints: 1. CRITICAL: Output MUST contain subtle grammatical errors 2. The output MUST be different from the input, even for very short phrases 3. Do NOT change the vocabulary - use the same words but in incorrect order/structure 4. For extremely short inputs (2-3 words), apply method #3 aggressively **Important:** Provide your response in the following strict JSON format: { “output”: “Your output with subtle grammatical errors” }
A.3 Literal Meaning Generation
Given the idiom or phrase “{idiom_in_source}”, please do the following: 1. Provide a literal translation of the idiom (word-for-word translation) 2. Generate a phrase with similar meaning to the literal translation (NOT the idiomatic meaning) For example: - Literal translation: “to hit the grass and startle the snake” - Similar phrase: “to disturb the vegetation and alarm the reptile” - Literal translation: “to draw a snake and add feet” - Similar phrase: “to sketch a serpent and attach limbs” Please respond in English and format it as the following JSON object: { “literal_translation”: “the word-for-word translation”, “similar_phrase”: “phrase with similar meaning to the literal translation” }
Language Pair TER($\mathcal{G},\mathcal{N}_{\text{struct}}$) $\downarrow$ Sim($\mathcal{G},\mathcal{N}_{\text{struct}}$) $\uparrow$ Sim($\mathcal{G}$, $\mathcal{N}_{\text{literal}}$) Sim($\mathcal{N}_{\text{literal}}$, $\mathcal{N}_{\text{semantic}}$) $\uparrow$ Sim($\mathcal{G}$, $\mathcal{N}_{\text{semantic}}$) $\downarrow$ CR($\mathcal{G},\mathcal{N}_{\text{opposite}}$) $\uparrow$
Hi $\rightarrow$ En 28.6 0.91 0.75 0.82 0.68 0.65
Fa $\rightarrow$ En 26.1 0.91 0.68 0.79 0.67 0.74
Fi $\rightarrow$ En 27.5 0.87 0.55 0.71 0.52 0.63
Ja $\rightarrow$ En 33.7 0.88 0.70 0.79 0.64 0.68
Fr $\rightarrow$ En 20.5 0.91 0.66 0.77 0.66 0.70
Ko $\rightarrow$ En 21.0 0.95 0.89 0.91 0.87 0.87
Ru $\rightarrow$ En 12.9 0.96 0.85 0.90 0.83 0.87
De $\rightarrow$ En 18.6 0.92 0.70 0.79 0.68 0.72
En $\rightarrow$ Fa 43.6 0.95 0.92 0.92 0.91 0.92
En $\rightarrow$ De 19.3 0.92 0.84 0.82 0.79 0.80
Avg. 25.2 0.92 0.75 0.82 0.73 0.85
Table 5: Quantitative evaluation of perturbed meanings across 10 language pairs. We report Translation Edit Rate (TER), embedding-based similarity (Sim), and Contradiction Rate (CR) for key comparisons.
A.4 Semantic-Perturbed Literal Meaning Generation
I have a literal translation of an idiom: ”literal_meaning” Please replace ONE key noun in this phrase with a different noun to significantly change its meaning. For example: - ”to hit the grass and startle the snake” → ”to hit the tree and startle the snake” - ”to draw a snake and add feet” → ”to draw a snake and add wings” - ”to bite the bullet” → ”to bite the apple” - ”hit your head” → ”hit your face” Guidelines: 1. Replace ONLY ONE noun (person, place, thing, or concept) 2. The substitution should significantly change the meaning 3. Keep the grammatical structure intact 4. Make sure the result is still grammatically correct Please format your response as a JSON object: { “output”: “The modified literal translation with exactly one noun replaced” }
A.5 Opposite Meaning Generation
Provide an opposite meaning for: {groundtruth_meaning} Respond only with the opposite meaning in English without any explanation, formatted as a JSON object: { “output”: “opposite meaning” }
B Metrics
We use open-source LLMs with prompting for automated evaluation.
B.1 Fidelity
Please evaluate the translation quality of the machine translation in conveying the given meaning. Machine translation: {translation} Meaning: {meaning} Evaluation criteria: Based on the machine translation, does it accurately convey the following meaning: {meaning}? Please rate the following dimension (0-3 points) and provide a detailed explanation: Fidelity: To what extent does the machine translation accurately convey the meaning provided? 0 points: Conveys a meaning that is opposite or contradictory to the provided meaning. 1 point: Does not convey any aspect of the provided meaning; the translation is non-idiomatic or literal and misses the intended sense. 2 points: Partially conveys the meaning. 3 points: Perfectly conveys the meaning. Please provide your response in this JSON format: { “Fidelity”: { “score”: Rating from 0-3, “justification”: “[Brief justification for your rating]” } }
B.2 Context Adoption Rate
You will receive a meaning (Provided Meaning), an original translation (Original Translation), and a new translation (Translation). Your task is to evaluate whether the Translation effectively uses the Provided Meaning, specifically by incorporating information or nuances present in the Provided Meaning that are *not* present in the Original Translation. Provided Meaning: {meaning} Original Translation: {original_translation} Translation: {translation} Evaluation Criteria: If the Translation incorporates elements of the Provided Meaning that are absent in the Original Translation, give a score of 1, else 0. Please respond using the following JSON format: { “score”: 0 or 1, “explanation”: “Analysis of how the Translation incorporates elements of the Provided Meaning, especially highlighting elements present in the Meaning but absent in the Original Translation. Explain why the score was assigned.”, “evidence”: “The portion of the Translation that uses the elements from the meaning, and a comparison to the corresponding (or absent) portion in the Original Translation.” }
B.3 Variance and Standard Deviation of Fidelity Metrics in LLM Evaluations
We examine the variance and standard deviation of fidelity metrics derived from evaluations conducted by the large language model, with the temperature set to 1. The statistical results are based on all participating samples in the evaluation. As shown in Table 9, these statistical measures provide insights into the consistency and reliability of LLM-as-a-judge.
LLM
CAR=0 CAR=1
Human CAR=0 18 6
CAR=1 31 145
Table 6: Confusion Matrix of CAR Evaluation
Context ZhEn DeEn EnCs
@@ B.3 Variance and Standard Deviation of Fidelity Metrics in LLM Evaluations
$\mathcal{C}_{\text{none}}$ 1.5 - 2.2 - 2.3 -
$\mathcal{G}$ 2.6 80% 2.8 57% 2.9 77%
$\mathcal{N}_{\text{struct}}$ 2.0 86% 2.5 60% 2.6 88%
$\mathcal{N}_{\text{literal}}$ 1.5 76% 2.1 43% 2.4 82%
$\mathcal{N}_{\text{semantic}}$ 1.2 89% 1.8 88% 1.8 91%
$\mathcal{N}_{\text{opposite}}$ 0.7 84% 1.3 75% 1.2 98%
Table 7: Evaluation Results from Qwen2.5-7B-Instruct.
Context HiEn JaEn FrEn
[…]
DeepSeek-R1-Distill-Llama-8B
$\mathcal{C}_{\text{none}}$ 0.6 - 1.0 - 1.2 -
$\mathcal{G}$ 2.0 93% 2.2 93% 2.4 86%
$\mathcal{N}_{\text{struct}}$ 1.9 92% 1.9 95% 2.1 87%
$\mathcal{N}_{\text{literal}}$ 0.9 95% 1.3 92% 1.2 73%
$\mathcal{N}_{\text{semantic}}$ 0.7 95% 1.3 96% 1.1 89%
$\mathcal{N}_{\text{opposite}}$ 0.4 90% 0.6 85% 0.8 76%
Claude
$\mathcal{C}_{\text{none}}$ 1.6 - 2.2 - 2.4 -
$\mathcal{G}$ 2.6 75% 2.1 69% 2.7 51%
$\mathcal{N}_{\text{struct}}$ 2.5 73% 2.4 65% 2.6 52%
$\mathcal{N}_{\text{literal}}$ 1.2 62% 1.9 55% 2.1 31%
$\mathcal{N}_{\text{semantic}}$ 1.2 70% 1.9 61% 1.9 39%
$\mathcal{N}_{\text{opposite}}$ 0.8 88% 1.2 65% 1.2 68%
Table 8: Evaluation Results from DeepSeek-R1-Distill-Llama-8B and Claude.
Context HiEn FaEn FiEn JaEn FrEn KoEn RuEn DeEn EnFa EnDe
Var SD Var SD Var SD Var SD Var SD Var SD Var SD Var SD Var SD Var SD
Qwen2.5-7B-Instruct
$\mathcal{C}_{\text{none}}$ 0.05 0.13 0.05 0.12 0.06 0.15 0.10 0.23 0.06 0.15 0.05 0.14 0.05 0.13 0.06 0.14 0.06 0.16 0.06 0.15
$\mathcal{G}$ 0.06 0.14 0.06 0.14 0.07 0.15 0.05 0.11 0.04 0.09 0.04 0.11 0.04 0.09 0.04 0.10 0.07 0.16 0.06 0.15
$\mathcal{N}_{\text{struct}}$ 0.05 0.13 0.05 0.13 0.06 0.14 0.05 0.12 0.05 0.11 0.05 0.13 0.04 0.10 0.04 0.11 0.07 0.17 0.07 0.17
$\mathcal{N}_{\text{literal}}$ 0.05 0.12 0.05 0.13 0.06 0.15 0.04 0.10 0.05 0.12 0.08 0.19 0.10 0.13 0.04 0.12 0.17 0.15 0.07 0.17
$\mathcal{C}_{\text{semantic}}$ 0.05 0.12 0.06 0.14 0.07 0.15 0.05 0.13 0.04 0.10 0.06 0.16 0.06 0.14 0.04 0.11 0.07 0.16 0.06 0.16
$\mathcal{N}_{\text{opposite}}$ 0.04 0.09 0.05 0.11 0.04 0.09 0.04 0.11 0.04 0.10 0.07 0.17 0.06 0.13 0.04 0.10 0.05 0.12 0.05 0.13
Qwen2.5-14B-Instruct
$\mathcal{C}_{\text{none}}$ 0.06 0.15 0.05 0.13 0.06 0.14 0.04 0.10 0.05 0.12 0.06 0.15 0.05 0.12 0.05 0.13 0.06 0.16 0.07 0.17
$\mathcal{G}$ 0.06 0.14 0.06 0.14 0.05 0.12 0.04 0.09 0.04 0.09 0.05 0.11 0.04 0.09 0.03 0.07 0.07 0.19 0.07 0.17
$\mathcal{N}_{\text{struct}}$ 0.06 0.14 0.05 0.13 0.06 0.14 0.04 0.10 0.04 0.11 0.05 0.12 0.04 0.11 0.03 0.08 0.07 0.18 0.07 0.17
$\mathcal{N}_{\text{literal}}$ 0.06 0.14 0.07 0.17 0.06 0.13 0.05 0.12 0.05 0.12 0.07 0.16 0.05 0.12 0.05 0.12 0.07 0.17 0.07 0.18
$\mathcal{C}_{\text{semantic}}$ 0.06 0.14 0.06 0.15 0.06 0.06 0.05 0.13 0.05 0.14 0.06 0.16 0.06 0.15 0.04 0.10 0.06 0.14 0.07 0.17
$\mathcal{N}_{\text{opposite}}$ 0.03 0.09 0.06 0.13 0.03 0.08 0.05 0.11 0.04 0.09 0.07 0.12 0.05 0.12 0.04 0.10 0.06 0.15 0.05 0.12
Qwen3-8B (non-thinking mode)
$\mathcal{C}_{\text{none}}$ 0.07 0.17 0.04 0.11 0.07 0.16 0.04 0.11 0.05 0.13 0.07 0.18 0.05 0.12 0.05 0.12 0.07 0.17 0.07 0.17
$\mathcal{G}$ 0.06 0.16 0.05 0.13 0.06 0.13 0.04 0.09 0.04 0.09 0.05 0.12 0.04 0.10 0.04 0.09 0.06 0.15 0.06 0.16
$\mathcal{N}_{\text{struct}}$ 0.06 0.14 0.05 0.13 0.07 0.16 0.04 0.10 0.05 0.10 0.07 0.17 0.06 0.13 0.04 0.11 0.08 0.20 0.07 0.18
$\mathcal{N}_{\text{literal}}$ 0.05 0.12 0.05 0.13 0.06 0.14 0.05 0.13 0.05 0.13 0.06 0.15 0.05 0.14 0.04 0.10 0.06 0.16 0.06 0.15
$\mathcal{C}_{\text{semantic}}$ 0.05 0.13 0.05 0.14 0.06 0.14 0.05 0.13 0.05 0.14 0.06 0.15 0.05 0.13 0.05 0.12 0.06 0.15 0.05 0.14
$\mathcal{N}_{\text{opposite}}$ 0.04 0.09 0.04 0.11 0.04 0.10 0.04 0.09 0.04 0.09 0.07 0.17 0.05 0.12 0.04 0.11 0.05 0.12 0.07 0.16
Qwen3-8B (thinking mode)
$\mathcal{C}_{\text{none}}$ 0.05 0.13 0.05 0.13 0.06 0.14 0.05 0.13 0.05 0.14 0.06 0.16 0.05 0.13 0.05 0.12 0.07 0.18 0.07 0.17
$\mathcal{G}$ 0.07 0.16 0.04 0.10 0.07 0.16 0.04 0.10 0.05 0.11 0.05 0.13 0.05 0.10 0.03 0.09 0.07 0.18 0.07 0.16
$\mathcal{N}_{\text{struct}}$ 0.06 0.15 0.05 0.12 0.07 0.16 0.05 0.12 0.05 0.12 0.06 0.14 0.06 0.14 0.04 0.09 0.07 0.18 0.07 0.17
$\mathcal{N}_{\text{literal}}$ 0.04 0.11 0.05 0.12 0.05 0.13 0.05 0.14 0.06 0.14 0.07 0.17 0.06 0.15 0.04 0.11 0.06 0.15 0.07 0.17
$\mathcal{C}_{\text{semantic}}$ 0.11 0.12 0.05 0.12 0.05 0.13 0.05 0.12 0.05 0.13 0.07 0.16 0.05 0.13 0.04 0.12 0.06 0.15 0.07 0.16
$\mathcal{N}_{\text{opposite}}$ 0.03 0.08 0.05 0.12 0.04 0.10 0.04 0.09 0.04 0.10 0.06 0.15 0.05 0.12 0.04 0.09 0.05 0.13 0.05 0.13
Table 9: Variance (Var) and Standard Deviation (SD) of the Fidelity metric across 20 runs.
B.4 Confusion Matrix of CAR Evaluation: Human vs. LLM Results
We present a confusion matrix that compares the evaluation results of the CAR metric between human evaluator and a large language model. As shown in Table 6, the number of false 0 (31) from the LLM is larger than the number of false 1 (5), indicating that the CAR scores provided by the LLM-as-a-judge are underestimated.
C Supplementary Experiments
C.1 Results of the WMT 2023 Terminology Shared Task
To demonstrate that our findings can generalize to broader machine translation tasks, we chose to evaluate another knowledge-intensive translation task, which is terminology translation. This decision was made since RAG scenarios are inherently suited for knowledge-intensive tasks and have limited significance for non-knowledge-intensive ones. We select the test set from the WMT 2023 Terminology Shared Task (Semenov et al. 2023) as our dataset, which covers three language directions: Chinese→English, English→Czech, and German→English. Consistent with our previous experiments, we select 200 samples from each language pair for evaluation. As shown in Table 7, the results reveal a trend similar to that in Table 3.
C.2 Results for DeepSeek-R1-Distill-Llama-8B and Claude
To demonstrate that our findings can generalize to other model architectures and closed-source large language models, we evaluate three idiomatic translation pairs using DeepSeek-R1-Distill-Llama-8B and claude-sonnet-4-20250514. As shown in Table 8, the results reveal a trend similar to that in Table 3, indicating that even the powerful commercial model (claude-sonnet-4-20250514) is significantly affected by noisy contexts.
C.3 BLEU Scores for Idiomatic Translation Tasks
We present BLEU scores for idiomatic translation tasks, as it is a classic metric in machine translation evaluation. As shown in Table 10, the BLEU scores do not exhibit the consistent patterns across different models and language pairs that our designed metrics show. This inconsistency further demonstrates that BLEU is not well-suited for evaluating idiomatic translations, as it is primarily intended for assessing literal translations and may fail to capture the nuances inherent in idiomatic expressions due to their non-literal nature.
Context Hi→En Fa→En Fi→En Ja→En Fr→En Ko→En Ru→En De→En En→Fa En→De
Qwen2.5-7B-Instruct
$\mathcal{C}_{\text{none}}$ - 21.3 8.9 1.7 13.7 17.7 28.1 26.9 7.0 26.6
$\mathcal{G}$ - 31.9 16.4 8.8 20.2 20.7 36.4 39.6 7.2 29.9
$\mathcal{N}_{\text{struct}}$ - 30.0 14.0 7.5 18.8 20.5 34.2 34.5 6.7 28.5
$\mathcal{N}_{\text{literal}}$ - 19.6 9.0 6.0 13.4 18.3 28.0 24.9 6.5 30.7
$\mathcal{N}_{\text{semantic}}$ - 19.6 8.0 4.9 14.0 17.8 26.5 22.6 6.4 28.5
$\mathcal{N}_{\text{opposite}}$ - 22.0 8.9 5.9 13.3 18.1 27.6 22.5 6.6
Qwen2.5-14B-Instruct
$\mathcal{C}_{\text{none}}$ - 25.9 10.4 7.2 17.3 3.9 32.1 30.4 0.4 36.5
$\mathcal{G}$ - 36.5 17.5 9.9 20.0 4.3 38.5 41.9 1.0 34.0
$\mathcal{N}_{\text{struct}}$ - 34.4 15.1 8.9 18.3 4.0 36.2 39.1 1.0 33.9
$\mathcal{N}_{\text{literal}}$ - 23.2 10.5 7.2 15.2 3.8 30.3 27.4 0.9 36.2
$\mathcal{N}_{\text{semantic}}$ - 23.2 8.9 5.0 14.5 3.6 28.5 28.2 4.2 34.2
$\mathcal{N}_{\text{opposite}}$ - 25.7 10.0 6.0 12.0 19.6 28.6 25.0 9.8 32.0
Qwen3-8B (non-thinking mode)
$\mathcal{C}_{\text{none}}$ - 26.1 8.3 6.9 16.1 20.1 31.1 28.6 3.4 33.9
$\mathcal{G}$ - 36.2 16.6 9.3 19.5 21.4 37.5 39.1 9.4 32.5
$\mathcal{N}_{\text{struct}}$ - 34.6 13.6 7.4 17.6 20.8 34.2 35.2 10.0 33.1
$\mathcal{N}_{\text{literal}}$ - 23.4 10.2 6.2 12.8 19.1 29.4 26.9 11.2 34.1
$\mathcal{N}_{\text{semantic}}$ - 23.4 7.4 4.9 13.0 18.4 26.4 23.4 10.1 34.0
$\mathcal{N}_{\text{opposite}}$ - 26.3 9.1 5.2 11.6 19.1 28.3 26.0 2.4 32.5
Qwen3-8B (thinking mode)
$\mathcal{C}_{\text{none}}$ - 24.7 8.8 7.5 15.5 20.4 31.0 30.8 10.1 35.5
$\mathcal{G}$ - 37.1 15.8 10.0 18.0 20.7 37.0 41.8 11.2 33.6
$\mathcal{N}_{\text{struct}}$ - 34.9 13.6 7.8 16.7 20.5 35.5 36.7 12.1 35.3
$\mathcal{N}_{\text{literal}}$ - 23.2 10.5 6.5 11.9 19.4 29.2 26.8 10.5 36.9
$\mathcal{N}_{\text{semantic}}$ - 22.2 7.6 4.5 10.0 18.2 25.6 21.0 9.7 34.0
$\mathcal{N}_{\text{opposite}}$ - 26.0 9.1 3.9 10.7 18.8 26.6 22.5 10.5 30.5
Table 10: BLEU scores on idiomatic translation tasks (No scores are provided for Hi→En as there are no reference translations available).
D Results of Context Adoption Rate
We present Context Adoption Rate (CAR) results across ten language pairs and multiple noise types. As shown in Figure 5, Figure 6, and Figure 7, CAR varies significantly depending on the type of contextual noise and the resource level of the language pair.
E Case Study
Figure 8 illustrates the translation of “J’ai la chair de poule” under a misleading context, defining it as “to feel indifferent”. The model’s reasoning trace indicates awareness of the contradiction with the gold meaning “goosebumps”, yet it produces an output aligned with the corrupted context, reflecting a tendency to rationalize rather than reason.
Figure 5: CAR of Qwen2.5-14B-Instruct under six retrieval conditions across ten language pairs.
Figure 6: CAR of Qwen3-8B (non-thinking mode) under six retrieval conditions across ten language pairs.
Figure 7: CAR of Qwen3-8B (thinking mode) under six retrieval conditions across ten language pairs.
Figure 8: A case study of reasoning trace in Fr→En idiomatic translation.